OpenStudy (anonymous):

What is the ordered pair for the vertex of the graph? f(x)=-2x^2+4x-6

OpenStudy (anonymous):

You want x=-b/2a. So here, a=-2 and b=4, so x=-4/2(-2)=1. Then calculate f(1) and that's your answer.

OpenStudy (anonymous):

In other words, the point you will plot is (1,f(1)).